# Simple Policy Optimization

### v5

@@ Abstract
PPO (Proximal Policy Optimization) algorithm has demonstrated excellent performance in many fields, and it is considered as a simple version of TRPO (Trust Region Policy Optimization) algorithm. However, the ratio clipping operation in PPO may not always effectively enforce the trust region constraints, this can be a potential factor affecting the stability of the algorithm. In this paper, we propose Simple Policy Optimization (SPO) algorithm, which introduces a novel clipping method for KL divergence between the old and current policies. Extensive experimental results in Atari 2600 environments indicate that, compared to the mainstream variants of PPO, SPO achieves better sample efficiency, extremely low KL divergence, and higher policy entropy, and is robust to the increase in network depth or complexity. More importantly, SPO maintains the simplicity of an unconstrained first-order algorithm. Our code is available at https://github.com/MyRepositories-hub/Simple-Policy-Optimization.
@@ 1 Introduction
Policy based model-free reinforcement learning algorithms have been widely applied in various fields, and have achieved performance surpassing humans including board games [1, 2, 3], video games [4, 5, 6], autonomous driving [7, 8] and intelligent control [9, 10, 11]. Although there are two main learning paradigms in deep reinforcement learning: value-based and policy-based, impressive algorithms such as PPO [12] and A3C [13] rely on the Actor-Critic (AC) framework [14], which is a synthesis of the two paradigms.
@@
Proximal Policy Optimization (PPO) [12] algorithm implicitly limits the difference between the old policy and the current policy through the ratio clipping operation, which is reflected in the Trust Region Policy Optimization (TRPO) [15] algorithm as a hard constraint on their KL divergence expectation. Ratio clipping avoids solving the expensive KL divergence constraint problem, so the PPO algorithm can also be well extended to large-scale reinforcement learning environments [4, 6]. Prior works have discussed how ratio clipping can effectively implement trust region constraints. For example, even if the probability ratio is bounded, the corresponding KL divergence is not necessarily bounded [16]. The performance improvement of PPO algorithm is likely to mainly depend on “code-level optimizations” [17]. There are also some works that design novel probability ratio clipping methods or objective functions [18, 19, 20, 21], or alternate between a standard RL phase and a distillation phase during the training stage to improve sample efficiency of PPO [22, 23].
There are many ways to avoid excessive KL divergence, which gives birth to many variants of PPO. For example, KL divergence can be added as a penalty term to the surrogate objective and introduce adaptive penalty coefficient [12]. During the training process of PPO, a KL threshold is set, and an early stopping strategy is adopted when the KL divergence reaches the threshold [24]. Alternatively, an early stopping strategy can be implemented based on the deviation degree of probability ratio instead of KL divergence [25].
@@
However, the above works may not have fundamentally changed the characteristics of the PPO algorithm (we acknowledge that the surrogate objective with ratio clipping remains the preferred choice in most cases), nor have they effectively constrained the KL divergence between the old and current policies. We propose Simple Policy Optimization (SPO) algorithm to compensate for these gaps, SPO does not even require the clipping of probability ratio, instead of a novel clipping method for the KL divergence between the old and current policies. Using SPO, we highlight three important observations about it:
@@
  • It can effectively limit the average KL divergence between the current policy and the old policy in almost all environments, while the PPO algorithm inevitably results in a high value of KL divergence between the old and current policies.
@@
  • It requires the calculation of the KL divergence between the old and current policies, which does not increase much computational overhead, and still, it maintains the simplicity of an unconstrained first-order algorithm.
@@
  • The aforementioned characteristics of the SPO algorithm are robust to the increase of network depth or complexity. In contrast, the increase in KL divergence during the training process of the PPO algorithm becomes more evident as the network deepens.
@@ 2.1 Policy Gradient
Reinforcement learning is generally defined by a tuple $(\mathcal{S},\mathcal{A},r,\mathcal{P},\rho_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state space and action space, $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the probability distribution of the state transition function, $\rho_{0}:\mathcal{S}\mapsto[0,1]$ is the initial state distribution, and $\gamma\in[0,1]$ is the discount factor.
@@
where $\theta$ is the parameters of the policy network, $R(\tau)=\sum_{i=1}^{T-1}\gamma^{i-1}r_{i}$, and the policy gradient can be estimated as
$\begin{split}\nabla J(\theta)&=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[R(\tau)\cdot\nabla\log p_{\theta}(\tau)\right]\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\log\left[\rho_{0}(s_{1})\cdot\prod_{t=1}^{T-1}\pi_{\theta}(a_{t}|s_{t})\cdot\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\left[\log\rho_{0}(s_{1})+\sum_{t=1}^{T-1}\log\pi_{\theta}(a_{t}|s_{t})+\log\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[\sum_{t=1}^{T-1}R(\tau)\cdot\nabla\log\pi_{\theta}(a_{t}|s_{t})\right]\approx\frac{1}{N}\sum_{n=1}^{N}\sum_{t=1}^{T_{n}-1}\left[R(\tau^{n})\cdot\nabla\log\pi_{\theta}(a_{t}^{n}|s_{t}^{n})\right],\\ \end{split}$ (3)
now we can estimate $\nabla J(\theta)$ through the Monte Carlo (MC) method [26].
@@ 2.2 Trust Region Policy Optimization
In TRPO, the following monotonic improvement theorem is given:
@@
[15] Let $\alpha=D_{\rm TV}^{\max}(\pi,\tilde{\pi})$ and $\epsilon=\max_{s}\lvert\mathbb{E}_{a\sim\tilde{\pi}}\left[A^{\pi}(s,a)\right]\rvert$, the following bound holds:
@@
$\begin{split}\max_{\theta}\enspace&\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t})\right]\\ \text{s.t.}\hskip 2.5pt\enspace&\mathbb{E}\left[D_{\text{KL}}(\pi_{\theta_{\rm old}},\pi_{\theta})\right]\leq\delta.\\ \end{split}$ (8)
@@ 2.3 Proximal Policy Optimization
PPO is an algorithm that attains the data efficiency and reliable performance of TRPO, while using only first-order optimization. PPO algorithm has two main variants, one of them is called PPO-Penalty, that is,
@@ 3 Method
In TRPO, trust region are imposed by limiting the expectation of the KL divergence between the old and current policies. Here, we denote the KL divergence under state $s_{t}$ as
@@
We will provide intuition about the objective (16), suppose that $d_{\max}=0.01$, and the function curves is shown in Figure 1. If $\hat{A}(s_{t},a_{t})>0$ and $d>d_{\max}$, the objective $J(\theta)$ will decrease rapidly (see curve $d_{\text{clip}}/d$), this can force the KL divergence to satisfy the trust region constraints. Otherwise, when $\hat{A}(s_{t},a_{t})<0$ and $d>d_{\max}$, the function curve $-d_{\text{clip}}/d+2$ and $d_{\text{clip}}/d$ are symmetric about the line $y=1$, and $J(\theta)$ will also decrease. Therefore, to maximize the objective $J(\theta)$, SPO will secretly optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ within the trust region $d\leq d_{\max}$ as much as possible, but it is still an unconstrained first-order algorithm.
@@ 4 Experiments
We preprocessed the Atari 2600 environment, for example, the input to the network is the stacking of the last four frames resized to $84\times 84\times 4$, reward clipping and so on [31, 32]. In all experiments, we use the hyperparameters provided in the Appendix A unless otherwise specified. For each environment in Atari 2600, we compute and visualize the standard deviation of the algorithm’s performance across 3 separate runs with different random seeds. Our experiments are divided into five parts:
@@ 4.2 Over Optimization
Figure 4: Shallow network and deeper network structure in Atari 2600 environments.
@@ 4.3 Increasing Network Complexity
We did not make any changes to the hyperparameters or other settings, except for increasing the number of network layers. PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks. The results are shown in Figure 5. As the network becomes deeper, we can see that the KL divergence of PPO during the training process becomes larger than that in the shallow networks, sometimes leading to a sharp decline in its performance. In some environments, PPO almost loses its ability to learn effectively (e.g., Breakout and DemonAttack). In contrast, SPO demonstrates strong robustness to changes in network structure. Furthermore, we observed that the average KL divergence curves of SPO-1 and SPO-2 nearly overlap throughout the entire training process, which once again validates the effectiveness of SPO in limiting KL divergence. Additionally, this suggests that SPO is more likely to play a pivotal role in complex, large-scale neural network models, such as the Transformers [34] or DMs (Diffusion Models) [35].
@@
Figure 6: Training curves in MuJoCo environments.
@@ 4.4 The Scalability of SPO for Continuous Action Space
The proof see Appendix B.6. For each environment in MuJoCo, we compute and visualize the standard deviation of the algorithm’s performance across 5 separate runs with different random seeds, the results are shown in Figure 6, Table 1 and Table 2.
Table 1: Average return of the last 10% training steps in MuJoCo environments.
@@
SPO-0.005 $4775.95^{\downarrow 440}$ $1623.82^{\color[rgb]{1,0,0}\uparrow 16}$ $1434.74^{\downarrow 37}$ $3587.51^{\color[rgb]{1,0,0}\uparrow 2765}$ $159441.18^{\color[rgb]{1,0,0}\uparrow 8996}$ $3187.8^{\downarrow 30}$
SPO-0.01 $5066.14^{\downarrow 150}$ $2328.6^{\color[rgb]{1,0,0}\uparrow 721}$ $1377.1^{\downarrow 95}$ ${\bf 3772.78}^{\color[rgb]{1,0,0}\uparrow 2950}$ $158407.45^{\color[rgb]{1,0,0}\uparrow 7962}$ ${\bf 3508.62}^{\color[rgb]{1,0,0}\uparrow 291}$
SPO-0.03 $4628.01^{\downarrow 588}$ ${\bf 2619.4}^{\color[rgb]{1,0,0}\uparrow 1012}$ ${\bf 1632.63}^{\color[rgb]{1,0,0}\uparrow 161}$ $3168.32^{\color[rgb]{1,0,0}\uparrow 2346}$ ${\bf 159841.75}^{\color[rgb]{1,0,0}\uparrow 9396}$ $3286.5^{\color[rgb]{1,0,0}\uparrow 69}$
PPO 5216.36 1607.55 1471.81 822.57 150445.3 3217.51
Table 2: Average return throughout the entire training process in MuJoCo environments.
@@
SPO-0.005 $1466.71^{\downarrow 524}$ $1229.05^{\downarrow 122}$ $1046.8^{\downarrow 109}$ $1318.69^{\color[rgb]{1,0,0}\uparrow 525}$ $132000.34^{\downarrow 4209}$ $1530.99^{\downarrow 362}$
SPO-0.01 $1249.85^{\downarrow 741}$ $1705.9^{\color[rgb]{1,0,0}\uparrow 355}$ $1088.38^{\downarrow 67}$ $1457.89^{\color[rgb]{1,0,0}\uparrow 664}$ $136563.02^{\color[rgb]{1,0,0}\uparrow 354}$ $1803.99^{\downarrow 89}$
SPO-0.03 $1035.62^{\downarrow 955}$ ${\bf 1861.74}^{\color[rgb]{1,0,0}\uparrow 511}$ $1132.27^{\downarrow 23}$ ${\bf 1501.02}^{\color[rgb]{1,0,0}\uparrow 708}$ ${\bf 136769.98}^{\color[rgb]{1,0,0}\uparrow 561}$ $1788.93^{\downarrow 104}$
PPO 1991.05 1351.08 1155.63 793.51 136209.27 1893.36
From Figure 6, we can see that the sample efficiency of SPO in the Humanoid-v4 environment is much better than PPO. This may be because the environment is relatively complex, and the KL divergence of PPO becomes too large (see Appendix B.6).
As we can see from Table 1, the SPO algorithm performs better than PPO in the last 10% training steps, especially for SPO-0.03, which outperforms PPO in all environments except Ant-v4. However, according to Table 2, the average performance of the SPO algorithm decreases throughout the entire training process. Overall, this indicates that SPO is capable of achieving or even surpassing the performance of PPO in continuous control tasks (especially tasks like Humanoid-v4, which are relatively more complex), because we value the final performance of the algorithm more than the average performance of the entire training process.

### v6

@@ Abstract
As one of the most important and influential algorithms in reinforcement learning, the Proximal Policy Optimization (PPO) algorithm has demonstrated outstanding performance across various domains. It simplifies the optimization-based importance sampling process of the Trust Region Policy Optimization (TRPO) algorithm through ratio clipping. However, this simplification with ratio clipping does not always effectively enforce trust region constraints. In this paper, we introduce an algorithm named Simple Policy Optimization (SPO), which incorporates a novel clipping method for the KL divergence between the old and new policies. Extensive experimental results in both Atari 2600 and MuJoCo environments show that, compared to PPO, SPO achieves better sample efficiency, extremely low KL divergence, and higher policy entropy, while also being robust to increases in network depth or complexity. More importantly, SPO maintains the simplicity of an unconstrained first-order algorithm. Our code is available at https://github.com/MyRepositories-hub/Simple-Policy-Optimization.
@@ 1 Introduction
Reinforcement learning has made rapid progress in recent years, becoming a cornerstone tool across a wide range of disciplines. There are two primary paradigms in deep reinforcement learning, value-based and policy-based. Policy-based model-free reinforcement learning algorithms have been extensively applied across various domains, achieving performances that surpass human capabilities in areas including board games [1, 2, 3], video games [4, 5, 6], autonomous driving [7, 8], and intelligent control systems [9, 10, 11]. Furthermore, notable algorithms like PPO [12] and A3C [13] utilize the Actor-Critic (AC) framework [14], which integrates these paradigms and serves as a foundational component in numerous fields, including games, robotics, large language models (LLMs), and more. In this paper, we focus on policy-based reinforcement learning algorithms and introduce the Simple Policy Optimization (SPO), which features a novel clipping method for importance sampling.
@@
The Proximal Policy Optimization (PPO) algorithm [12] limits the difference between the old and new policies through a ratio clipping operation, which in the Trust Region Policy Optimization (TRPO) algorithm [15] is manifested as a hard constraint on expected KL divergence. Ratio clipping circumvents the need to solve the computationally expensive KL divergence constraint problem, thus allowing the PPO algorithm to be effectively extended to large-scale reinforcement learning environments [4, 6]. Previous studies have explored the relationship between ratio clipping and trust region constraints. For instance, even if the probability ratio is bounded, the corresponding KL divergence does not necessarily be bounded [16]. The performance improvements in the PPO algorithm are likely primarily dependent on “code-level optimizations” [17]. Additionally, some works have designed novel probability ratio clipping methods or objective functions [18, 19, 20, 21], or alternated between a standard RL phase and a distillation phase during the training stage to enhance the sample efficiency of PPO [22, 23].
There are numerous strategies to mitigate excessive KL divergence, leading to various adaptations of the PPO algorithm. For instance, KL divergence can be incorporated as a penalty term within the surrogate objective, complemented by an adaptive penalty coefficient [12]. During PPO’s training phase, a predefined KL threshold is established. An early stopping mechanism is triggered once this KL divergence threshold is reached [24]. Alternatively, the early stopping strategy can be based on the deviation in the probability ratio rather than on KL divergence itself [25].
@@
However, the aforementioned approaches have not fundamentally altered the characteristics of the PPO algorithm (we acknowledge that the surrogate objective with ratio clipping remains the preferred choice in most cases), nor have they effectively constrained the KL divergence between the old and new policies. We propose the Simple Policy Optimization (SPO) algorithm to address these shortcomings. Notably, SPO does not require the clipping of the probability ratio; instead, it introduces a novel clipping method for the KL divergence between the old and new policies. We highlight three important contributions in SPO:
@@
  • SPO can effectively limit the average KL divergence between the old and new policies in almost all environments, while the PPO algorithm inevitably results in a high value of KL divergence between the successive policies.
@@
  • SPO requires the calculation of the KL divergence between successive policies, which does not increase much computational overhead, and still, it maintains the simplicity of an unconstrained first-order algorithm.
@@
  • The aforementioned characteristics of SPO are robust to increases in network depth or complexity. In contrast, the increases in KL divergence between successive policies of PPO becomes more evident as the network deepens.
@@ 2.1 Policy Gradient
Reinforcement learning is generally defined by a tuple $(\mathcal{S},\mathcal{A},r,\mathcal{P},\rho_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state space and action space, $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the probability distribution of the state transition function, $\rho_{0}:\mathcal{S}\mapsto[0,1]$ is the initial state distribution, and $\gamma\in(0,1]$ is the discount factor.
@@
where $\theta$ is the parameters of the policy network, $R(\tau)=\sum_{t=0}^{T-1}\gamma^{t}r_{t}$. Then the policy gradient can be estimated as
$\begin{split}\nabla J(\theta)&=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[R(\tau)\cdot\nabla\log p_{\theta}(\tau)\right]\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\log\left[\rho_{0}(s_{0})\cdot\prod_{t=0}^{T-1}\pi_{\theta}(a_{t}|s_{t})\cdot\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\left[\log\rho_{0}(s_{0})+\sum_{t=0}^{T-1}\log\pi_{\theta}(a_{t}|s_{t})+\log\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[\sum_{t=0}^{T-1}R(\tau)\cdot\nabla\log\pi_{\theta}(a_{t}|s_{t})\right]\approx\frac{1}{N}\sum_{n=1}^{N}\sum_{t=0}^{T_{n}-1}\left[R(\tau^{n})\cdot\nabla\log\pi_{\theta}(a_{t}^{n}|s_{t}^{n})\right],\\ \end{split}$ (3)
and now we can estimate $\nabla J(\theta)$ through the Monte Carlo (MC) method [26].
@@ 2.2 Trust Region Policy Optimization
In TRPO [15], the following monotonic improvement theorem is given:
@@
Let $\alpha=D_{\rm TV}^{\max}(\pi,\tilde{\pi})$ and $\epsilon=\max_{s,a}\left|A_{\pi}(s,a)\right|$, the following bound holds:
@@
$\begin{split}\max_{\theta}\enspace&\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t})\right],\\ \text{s.t.}\hskip 2.5pt\enspace&\mathbb{E}\left[D_{\text{KL}}(\pi_{\theta_{\rm old}},\pi_{\theta})\right]\leq\delta.\\ \end{split}$ (8)
@@ 2.3 Proximal Policy Optimization
PPO [12] is an algorithm that attains the data efficiency and reliable performance of TRPO, while using only first-order optimization. PPO algorithm has two main variants, one of them is called PPO-Penalty, that is,
@@ 3 Method
In TRPO, trust region are imposed by limiting the expectation of the KL divergence between the old and new policies. Here, we denote the KL divergence under state $s_{t}$ as
@@
We will provide intuition about the objective (16), suppose that $d_{\max}=0.01$, and the function curves are shown in Figure 1. If $\hat{A}(s_{t},a_{t})>0$ and $d>d_{\max}$, the objective $J(\theta)$ will decrease rapidly (see curve $d_{\text{clip}}/d$), this can force the KL divergence to satisfy the trust region constraints. Otherwise, when $\hat{A}(s_{t},a_{t})<0$ and $d>d_{\max}$, the function curve $-d_{\text{clip}}/d+2$ and $d_{\text{clip}}/d$ are symmetric about the line $y=1$, and $J(\theta)$ will also decrease. Therefore, to maximize the objective $J(\theta)$, SPO will secretly optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ within the trust region $d\leq d_{\max}$ as much as possible, but it is still an unconstrained first-order algorithm.
@@ 4 Experiments
We preprocessed the Atari 2600 environment, for example, the input to the network is the stacking of the last four frames resized to $84\times 84\times 4$, reward clipping and so on [31, 32]. In all experiments, we use the hyperparameters provided in the Appendix A unless otherwise specified. For each environment in Atari 2600, we compute and visualize the standard deviation of the algorithm’s performance across three separate runs with different random seeds. Our experiments are divided into five parts:
@@ 4.2 Over Optimization
Figure 4: Shallow network and deeper network structures in Atari 2600 environments.
@@ 4.3 Increasing Network Complexity
We did not make any changes to the hyperparameters or other settings, except for increasing the number of network layers. PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks. The results are shown in Figure 5. As the network becomes deeper, we can see that the KL divergence of PPO during the training process becomes larger than that in the shallow networks, sometimes leading to a sharp decline in its performance. In some environments, PPO almost loses its ability to learn effectively (e.g., Breakout and DemonAttack). In contrast, SPO demonstrates strong robustness to changes in network structure. Furthermore, we observed that the average KL divergence curves of SPO-1 and SPO-2 nearly overlap throughout the entire training process, which once again validates the effectiveness of SPO in limiting KL divergence. Additionally, this suggests that SPO is more likely to play a pivotal role in complex, large-scale neural network models, such as the Transformer [34] or DMs (Diffusion Models) [35].
@@
Figure 6: Training curves in MuJoCo with deeper networks.
@@ 4.4 The Scalability of SPO for Continuous Action Space
The details see Appendix B.7. For each environment in MuJoCo, we compute and visualize the standard deviation of the algorithm’s performance across five separate runs with different random seeds, and we also employ two different network architectures with varying depths (see Figure 16). The results are shown in Figure 6, Table 1, Table 2 and Figure 7.
Table 1: Average return of the last 10% training steps with shallow networks in MuJoCo.
@@
PPO 5216.4 1607.5 1471.8 822.6 150445.3 3217.5
SPO 4875.4 2459.9 1402.6 3228.6 163008.1 3520.9
Table 2: Average return of the last 10% training steps with deeper networks in MuJoCo.
@@
PPO 527.4 2781.0 1182.6 576.7 136717.4 2512.4
SPO 2679.8 3276.0 1390.8 3231.8 152980.9 3800.0
Figure 7: Error bar chart (from left to right): PPO, SPO-0.01, SPO-0.02, and SPO-0.03 with deeper networks.
Figure 6 and Table 2 clearly show that SPO outperforms PPO in sample efficiency across all environments when training with deeper neural networks. Furthermore, as can be seen from Figure 7, the KL divergence of SPO remains controlled during the training process, whereas the KL divergence of PPO becomes significantly large, particularly in Ant-v4 and Humanoid-v4 (the gray line in Figure 7 corresponds to a value of 0.02). This divergence led to the collapse of PPO’s performance in these environments.
Due to the limitation of space, please refer to Appendix B.7 for more detailed experimental results. In summary, we speculate that the performance of the PPO algorithm mainly depends on two factors: “code-level optimizations” and a relatively simple network structure. This is why, in commonly used benchmarks, the policy and value networks of PPO are always designed to be relatively simple. While this avoids excessive KL divergence, it also limits the algorithm’s capacity to encode complex tasks.